# Searching for periodic signals and quasi-periodic oscillations from an extremely active cycle of FRB 20240114A

### v1

@@ 4.3 Periodicity and QPOs in Burst Time Series
We detected QPOs only in a subset of radio pulses, likely because substructure detectability depends critically on time resolution and the signal-to-noise ratio, combined with possible source intensity variations. Even under similar observational conditions, quasi-periodic subpulses are not universally present. For example, the magnetar PSR J1622–4950 clearly exhibited such structures in 2017 observations, although they were relatively weak in observations from early and late 2018 (Kramer et al., 2024).
5 Summary
This study presents a comprehensive multi-timescale analysis of periodic signals and QPOs in the repeating fast radio burst FRB 20240114A, based on extensive observations with the FAST. Our investigation reveals several key findings across different temporal domains:
  1. 1.
    Short-Timescale Periodicity Search: We analyzed 11,553 bursts from 57 observations spanning January to August 2024. We identified three candidate periodic signals with periods of 0.673 s, 0.635 s, and 0.536 s (significance $>3\sigma$), each detected in two independent observations. It is clear that not all of these signals are related to the periodicity of the source. Nevertheless, our current selection criteria are not sufficient to definitively rule out any of them. The significant variation in the $\dot{P}$ across different observations is unlikely to be caused by the intrinsic spin-down of a magnetar. This suggests that either the candidates are due to stochastic processes, or the FRB source resides in a binary system, where the apparent period derivative variations are due to Doppler modulation. Despite the large sample size and the detected short-timescale periodic signals, a definitive period could not be confirmed, indicating that the link between the radiation characteristics of the FRB and spin periodicity may be weak.
@@ 5 Summary
    Long-Timescale Periodicity Analysis: Through epoch folding ($3.9\sigma$) and Lomb-Scargle periodogram ($3.5\sigma$) techniques, we detected a potential $\sim$68-day activity cycle. While the significance is limited by uneven observational coverage, if this periodicity is confirmed, it would place FRB 20240114A among the growing class of periodically active repeaters, possibly indicating binary orbital modulation, precessional effects, or other factors. The lack of robust detection highlights the need for more uniform future observations to constrain long-timescale periodicity in repeating FRBs.
@@
    Periodicity of burst clusters and QPOs within bursts: We conducted a comprehensive periodicity analysis on the time series of individual bursts and burst clusters to investigate both periodic signals and QPOs in FRBs. For the bursts analysis, our approach, which combines Bayesian inference with EMD, identified 5 significant QPOs ($>3\sigma$), with frequencies in the few hundred Hz range. These QPOs show a strong correlation between central frequency and width (Pearson $R=0.97$), suggesting that the mechanism behind the QPOs may be relatively stable. For the burst clusters analysis, autocorrelation function analysis of burst clusters identified several significant periodic candidates, ranging from a few milliseconds to several tens of milliseconds, with transient coherence over approximately 3–4 cycles. This suggests that oscillation modes are temporarily excited between bursts, providing insight into the magnetospheric plasma dynamics. Collectively, these findings highlight the presence of organized structures both within and between bursts, offering valuable constraints on the physical conditions of the emission region. Based on the relationship between the timescale of pulse microstructure and the rotational period of the source as given in Kramer et al. (2024), we suggest that the possible rotational period of the FRB source is between 2 s and 76 s, although we did not find this period in the data.
FRB 20240114A is currently one of the most active known repeating fast radio bursts, and the results of our periodicity analysis offer important statistical constraints on its physical properties. Through multi-timescale searches for periodic signals and QPOs, we have provided observational clues that may shed light on the source’s rotational characteristics, precession behavior, and dynamic processes occurring both within and between bursts. While these preliminary findings offer a useful starting point, these results clarify the bias of previous studies that detected occasional QPOs believed to be potentially related to the spin of compact objects. As data coverage expands, we hope to refine our understanding of the periodicity and enhance our insights into the connection between fast radio bursts, their rotational properties, and potential precession effects.
acknowledgments
This work made use of the data from FAST FRB Key Science Project. This work is supported by the National Natural Science Foundation of China (NSFC) under grant numbers 12588202, 12303042, 12041303, 12421003, 12233002, 12041306, 12403100, W2442001, 12203045, 12447115, U2031117 and the National SKA Program of China (2020SKA0120100, 2020SKA0120200, 2022SKA0130104). Di Li is a New Cornerstone investigator. Pei Wang acknowledges support from the Youth Innovation Promotion Association CAS (id. 2021055), CAS Youth Interdisciplinary Team. Di Li is supported by the International Partnership Program of Chinese Academy of Sciences, Program No.114A11KYSB20210010, National Key R&D Program of China No. 2023YFE0110500, QN2023061004L. Yi Feng is supported by the Leading Innovation and Entrepreneurship Team of Zhejiang Province of China grant No. 2023R01008, and by Key R&D Program of Zhejiang grant No. 2024SSYS0012. Yongfeng Huang also acknowledges the support from the Xinjiang Tianchi Program. Qin Wu is supported by the China Postdoctoral Science Foundation (CPSF) under grant numbers GZB20240308, 2025T180875. C.W.T is supported by CAS project No. JZHKYPT-2021-06. This research is also supported by the CAS Project for Young Scientists in Basic Research, YSBR-063.
Appendix A Search Method for Periodic Signals and QPOs in bursts
@@
To reliably identify periodic signals and quasi-periodic oscillations (QPOs) in the bursts, we adopt the method proposed by Huppenkothen et al. (2013), which is based on Bayesian inference. The procedure begins by computing the periodogram of the burst time series using the Fast Fourier Transform (FFT). A simpler model, referred to as the null hypothesis model, is then selected, and the periodogram is fit using maximum likelihood estimation (MLE) to obtain the a maximum posteriori (MAP) parameters under this null hypothesis. The analysis is then performed separately for searching periodic signals and QPOs.
@@
$p\left(\bm{I}|\hat{\theta}_{\rm MAP}^{0},H_{0}\right)=\prod\limits_{i=1}^{n}p\left(I_{j}|S_{j}\right),$ (A2)
where $p(I_{j}|S_{j})$ represents the probability density function (PDF) of the periodogram at frequency $f_{j}$ given the model $S_{j}$. The posterior distribution of the null hypothesis model parameters is computed by combining this likelihood with the prior distribution. Markov Chain Monte Carlo (MCMC) sampling is used to obtain a large number of simulated periodograms based on the posterior distribution. The simulated periodograms are then fit with the null hypothesis model, and the maximum residuals for each simulated periodogram are calculated. The $P$-value is determined by comparing the maximum residual observed in the actual burst periodogram with those from the simulated ones. If the $P$-value is below a predefined significance level, the outlier is considered statistically significant and warrants further investigation.
@@
$T_{\rm LRT}^{\rm obs}=-2\log\frac{p\left(\bm{I}|\hat{\theta}_{\rm MAP}^{0},H_{0}\right)}{p\left(\bm{I}|\hat{\theta}_{\rm MAP}^{1},H_{1}\right)}.$ (A3)
@@
The PDF at each frequency in the periodogram is crucial for constructing the likelihood. Many researchers assume that the signal’s power spectral model is multiplied by white noise (Timmer & König, 1995; Vaughan, 2010; Huppenkothen et al., 2013; Li et al., 2022; Xiao et al., 2022), which conflicts with the use of multiple additive models to fit the periodogram. To resolve this issue, we propose a new periodogram PDF based on a time-domain additive model, which more accurately reflects the superposition of multiple components. This approach aligns with Guidorzi (2011), who reached similar conclusions. We first review the periodogram PDF under the frequency-domain multiplication model assumption and then present the PDF under the time-domain additive model assumption.
In the frequency-domain multiplication model, consider a Gaussian white noise $w(t)$ with N time samples, mean $\mu$, and standard deviation $\sigma$. As is well known, its power spectrum at $f>0$ follows ${\rm PDS}(f)=\sigma^{2}$. The periodogram is used to estimate the power spectrum, which can be expressed as
${\rm Per_{W}}(f)=|W(f)|^{2}=\left|\frac{1}{N}\sum_{t=0}^{N-1}w(t)e^{-j2\pi ft}\right|^{2},$ (A4)
where $W(f)$ is the Fourier transform of $w(t)$. It is known that ${\rm Per_{w}}(f)=\frac{1}{2}\sigma^{2}\chi_{2}^{2}$ (Timmer & König, 1995), where $\chi_{2}^{2}$ is a random variable following a chi-squared distribution with 2 degrees of freedom. Therefore, the PDF for the power value ${\rm Per_{w}}(f)=y\ (y>0)$ is
${\rm PDF_{W}}(y;f)=\frac{1}{\sigma^{2}}e^{-\frac{y}{\sigma^{2}}},\quad y>0.$ (A5)
This expression is also equivalent to an exponential distribution with parameter $\lambda=\frac{1}{\sigma^{2}}$. It is considered that for linear processes, the total power spectrum is obtained by multiplying the power spectrum ${\rm PDS_{s}}(f)$ representing the process with the white noise power spectrum ${\rm PDS_{w}}(f)$ (Timmer & König, 1995), i.e., the total power spectrum is ${\rm PDS_{M}}(f)={\rm PDS_{s}}(f)\times{\rm PDS_{w}}(f)$. Therefore, for $f>0$, the probability density of the total periodogram ${\rm Per}(f)=y\ (y>0)$ is
${\rm PDF_{M}}(y;f)=\frac{1}{{\rm PDS}(f)}e^{-\frac{y}{{\rm PDS}(f)}},\quad y>0.$ (A6)
This is consistent with the methods used by Vaughan (2010), Li et al. (2022) and Xiao et al. (2022).
In the time-domain addition model, consider a signal $x(t)$ with N time samples, which consists of a deterministic pulse signal $s(t)$ and a Gaussian white noise signal $w(t)$ with mean $\mu$ and standard deviation $\sigma$, i.e.:
$x(t)=s(t)+w(t).$ (A7)
The Fourie transform of the signal $x(t)$ is
$X(f)=S(f)+W(f),$ (A8)
where $S(f)$ is the Fourie transform of the signal $s(t)$, and the value is deterministic. $W(f)$ is the Fourie transform of the noise $w(t)$, and the value is a complex Gaussian random variable (Timmer & König, 1995), i.e., $W(f)=\mathcal{N}(0,\frac{1}{2}\sigma^{2})+j\mathcal{N}(0,\frac{1}{2}\sigma^{2})$. The total periodogram ${\rm Per_{A}}(f)\equiv|X(f)|^{2}$, and by substituting $X(f)=S(f)+W(f)$, we get
${\rm Per_{A}}(f)=\left|S(f)+W(f)\right|^{2}=\left[{\rm Re}(S(f))+{\rm Re}(W(f))\right]^{2}+\left[{\rm Imag}(S(f))+{\rm Imag}(W(f))\right]^{2}.$ (A9)
It is clear that ${\rm Re}(S(f))+{\rm Re}(W(f))\sim\mathcal{N}({\rm Re}(S(f)),\frac{1}{2}\sigma^{2})$, and ${\rm Imag}(S(f))+{\rm Imag}(W(f))\sim\mathcal{N}({\rm Imag}(S(f)),\frac{1}{2}\sigma^{2})$. Therefore, the power value ${\rm Per_{A}}(f)=\frac{\sigma^{2}}{2}\chi_{2}^{2}\left(\frac{2|S(f)|^{2}}{\sigma^{2}}\right)$, where $\chi_{2}^{2}\left(\frac{2|S(f)|^{2}}{\sigma^{2}}\right)$ is a non-central chi-squared random variable with 2 degrees of freedom and the non-centrality parameter $\lambda(f)=\frac{2|S(f)|^{2}}{\sigma^{2}}$. The PDF for the power value ${\rm Per}(f)=y\ (y>0)$ is
${\rm PDF_{A}}(y;f)=\frac{1}{\sigma^{2}}e^{-\frac{y+|S(f)|^{2}}{\sigma^{2}}}I_{0}\left(\sqrt{\frac{4y|S(f)|^{2}}{\sigma^{4}}}\right),\quad y>0,$ (A10)
where $I_{0}(z)=\sum_{n=0}^{\infty}\frac{1}{(n!)^{2}}\left(\frac{z}{2}\right)^{2n}$ is the modified Bessel function of the first kind, zeroth order. Note that when $|S(f)|^{2}=0$ in equation (A10), the equation degenerates into the result of pure white noise, as shown in equation (A5).
@@
Our search strategy follows that of Huppenkothen et al. (2013) with the following key adjustments: the null hypothesis model consists of a Gaussian model and a constant function model. The Gaussian model fits the low-frequency component of the periodogram, while the constant function model fits the baseline of the periodogram. We found that the Gaussian model provides the best fit for the low-frequency component, and additional Gaussian models are added to fit narrower features in the periodogram. This results in a nested model that serves as the alternative hypothesis. When constructing the likelihood and simulating the periodogram, we use the time-domain additive model PDF (equation A10) rather than the frequency-domain multiplication model PDF (equation A6). After the initial candidate is identified, the periodogram is reviewed manually to determine if additional Gaussian models are needed. If the periodogram cannot be fit with at most five Gaussian functions plus one constant function (with a reduced $\chi^{2}>1.5$), the candidate is discarded.
@@
Appendix B QPO fitting parameters for individual bursts
Table 2: QPO fitting parameters for individual bursts. The table lists the model components (Model), their associated amplitudes (Height), central frequencies (CF), full width at half maximum (FWHM), quality factors (Q), and constant (C) for the QPO in each individual burst. The MJD in each row indicates the start time of the time series for that particular burst.
No. MJD DM Model Height CF FWHM Q C
From 60355 $\rm pc/cm^{3}$ Hz Hz ×$\times 10^{-3}$
1 20.182604131 528.5 G0 6.337 $\pm$ 0.046 4.45 $\pm$ 0.59 85.02 $\pm$ 0.63 0.0523 $\pm$ 0.0069 -
G1 0.1865 $\pm$ 0.0037 266.85 $\pm$ 0.68 97.0 $\pm$ 1.3 2.750 $\pm$ 0.037 -
G2 0.0053 $\pm$ 0.0011 502.1 $\pm$ 6.2 70 $\pm$ 14 7.1 $\pm$ 1.4 -
Const - - - - $3.0\pm 4.6$
2 20.200749305 527.8 G0 160.4 $\pm$ 2.3 0.8 $\pm$ 1.4 115.0 $\pm$ 1.4 0.007 $\pm$ 0.012 -
G1 14.22 $\pm$ 0.18 277.90 $\pm$ 0.61 124.7 $\pm$ 1.3 2.228 $\pm$ 0.024 -
G2 0.592 $\pm$ 0.029 501.8 $\pm$ 8.3 212 $\pm$ 12 2.36 $\pm$ 0.14 -
Const - - - - $2.7\pm 2.9$
3 21.143877879 528.7 G0 36.3 $\pm$ 0.4 5.8 $\pm$ 1.1 118.5 $\pm$ 1.1 0.0487 $\pm$ 0.0094 -
G1 0.946 $\pm$ 0.021 327.94 $\pm$ 0.93 117.4 $\pm$ 1.7 2.793 $\pm$ 0.041 -
G2 0.0368 $\pm$ 0.0047 579.8 $\pm$ 5.6 105 $\pm$ 10 5.50 $\pm$ 0.55 -
Const - - - - $2.2\pm 1.2$
4 26.127338551 527.3 G0 9.04 $\pm$ 0.28 0.0 $\pm$ 2.7 253.2 $\pm$ 5.9 $0$$\pm$ 0.011 -
G1 0.212 $\pm$ 0.012 586.6 $\pm$ 4.4 212.0 $\pm$ 8.6 2.77 $\pm$ 0.11 -
Const - - - - $2.2\pm 2.0$
5 114.836935924 528.3 G0 6.02 $\pm$ 0.14 0.0 $\pm$ 2.9 176.0 $\pm$ 2.7 $0$$\pm$ 0.016 -
G1 0.0980 $\pm$ 0.0062 468.1 $\pm$ 3.9 159.8 $\pm$ 7.1 2.93 $\pm$ 0.13 -
Const - - - - $2.3\pm 1.4$
Appendix C Quasi-Periodic Signals in Burst Time Series
Figure 9: Detections of QPOs in the burst time series of FRB 20240114A. Each candidate is displayed in a panel of four subplots, with its identifier labeled in the top-left subplot. For each candidate, the left column shows the original burst time series, the EMD result (black curve), and the fitted model from equation (8) (pink curve), while the right column displays the corresponding periodogram, where dashed lines indicate the best-fit components and the central frequencies of QPOs.

### v2

@@ 5 Summary
    Using FAST observations with TOAs spanning from UT 2024-01-28 to UT 2025-10-01, we detected a significant long-timescale periodicity at $143.40\pm 7.19$ days with $5.2\sigma$ significance. The coherent phase folding over the entire $\sim$20-month observational baseline demonstrates consistent burst concentration in specific phase regions, establishing FRB 20240114A as a key member of the periodic repeater population.
@@
[…]
The authors declare that they have no conflict of interest.
Appendix A Performance of the TDA Pipeline on Simulated TOA Data
To evaluate the Time-Differencing Algorithm (TDA) in short-timescale periodicity searches, we simulated 300 TOAs over 1500 s with injected parameters $f_{0}=3.1415~\mathrm{Hz}$ and $f_{1}=1.0\times 10^{-6}~\mathrm{Hz/s}$, resembling the observational characteristics of FRB 20240114A. The analysis followed the main text pipeline: a TDA pre-screening over $f_{1}/f_{0}$ to locate candidate periodicities, followed by refinement using the $Z_{n}^{2}$ statistic.
[…]
Using this pipeline, the most significant TDA peak was found at $f_{1}/f_{0}=3.1679\times 10^{-7}\rm s^{-1}$, corresponding to an outlier near $f_{0}=3.1417~\mathrm{Hz}$ (Figure 6, top left). From this $f_{0}$, the estimated frequency derivative is $f_{1}=9.9525\times 10^{-7}~\mathrm{Hz/s}$, closely matching the injected values. Fitting the complementary cumulative distribution function (CCDF) of the TDA spectrum with an exponential model (top right) yields a false-alarm probability of $2.18\times 10^{-10}$, corresponding to a $6.34\sigma$ significance.
[…]
These simulations demonstrate that TDA effectively identifies candidate periodicities in sparse TOA datasets. When combined with $Z_{n}^{2}$ refinement, the hybrid approach achieves both computational efficiency and accurate recovery of injected periods, supporting its application for short-timescale periodicity searches in FRB-like datasets.
Appendix B FAST Observational Log of FRB 20240114A
Table 3: FAST observational log of FRB 20240114A used in the long-timescale periodicity analysis. The table lists observation dates, MJD for start and end times, observation durations in hours, number of detected bursts, and burst rates in bursts per hour.
Date (UT) Start MJD End MJD Duration (hr) Burst Count Rate (hr-1)
20240128 60337.229861 60337.250694 0.50 8 16.0
20240129 60338.215278 60338.236111 0.50 11 22.0
20240201 60341.179861 60341.200694 0.50 9 18.0
20240204 60344.240972 60344.261806 0.50 2 4.0
20240209 60349.152778 60349.173611 0.50 6 12.0
20240213 60353.215972 60353.236806 0.50 14 28.0
20240215 60355.131250 60355.152083 0.50 21 42.0
20240221 60361.179167 60361.200000 0.50 24 48.0
20240225 60365.131944 60365.152778 0.50 11 22.0
20240305 60374.188194 60374.209028 0.50 225 450.0
20240306 60375.134722 60375.218056 2.00 918 459.0
20240307 60376.102778 60376.186111 2.00 969 484.5
20240308 60377.193056 60377.221458 0.68 347 509.0
20240309 60378.191667 60378.205671 0.34 186 553.4
20240310 60379.199306 60379.219132 0.48 270 567.4
20240311 60380.152778 60380.194444 1.00 727 727.0
20240312 60381.038194 60381.220833 4.38 3197 729.4
20240313 60382.173611 60382.215278 1.00 535 535.0
20240315 60384.190972 60384.204861 0.33 66 198.0
20240317 60386.118750 60386.139583 0.50 171 342.0
20240321 60390.158333 60390.172222 0.33 111 333.0
20240327 60396.133333 60396.154167 0.50 297 594.0
20240329 60398.981944 60399.002778 0.50 275 550.0
20240401 60401.066667 60401.087500 0.50 261 522.0
20240403 60403.046528 60403.067361 0.50 141 282.0
20240407 60407.091667 60407.112500 0.50 41 82.0
20240410 60410.020833 60410.041667 0.50 35 70.0
20240413 60413.023611 60413.044444 0.50 44 88.0
20240417 60417.078472 60417.099306 0.50 45 90.0
20240421 60421.009028 60421.029861 0.50 72 144.0
20240424 60424.983333 60425.004167 0.50 110 220.0
20240430 60430.029167 60430.050000 0.50 133 266.0
20240504 60434.956944 60434.977778 0.50 89 178.0
20240511 60441.888889 60441.909722 0.50 102 204.0
20240518 60448.897222 60448.911111 0.33 32 96.0
20240525 60455.913889 60455.927778 0.33 37 111.0
20240531 60461.949306 60461.963194 0.33 11 33.0
20240608 60469.827778 60469.841667 0.33 49 147.0
20240615 60476.877083 60476.890972 0.33 29 87.0
20240622 60483.881944 60483.895833 0.33 40 120.0
20240629 60490.862500 60490.876389 0.33 46 138.0
20240706 60497.858333 60497.872222 0.33 42 126.0
20240713 60504.761111 60504.775000 0.33 117 351.0
20240720 60511.684722 60511.698611 0.33 100 300.0
20240722 60513.800694 60513.815278 0.35 92 262.9
20240724 60515.773611 60515.834722 1.47 369 251.6
20240726 60517.819444 60517.840278 0.50 166 332.0
20240727 60518.742361 60518.756250 0.33 92 276.0
20240728 60519.765278 60519.786111 0.50 208 416.0
20240730 60521.800000 60521.820833 0.50 138 276.0
20240802 60524.783333 60524.797222 0.33 108 324.0
20240805 60527.777778 60527.791667 0.33 52 156.0
20240808 60530.775000 60530.788889 0.33 90 270.0
20240814 60536.777778 60536.791667 0.33 86 258.0
20240817 60539.709722 60539.723611 0.33 70 210.0
20240821 60543.737500 60543.751389 0.33 66 198.0
20240829 60551.699306 60551.713194 0.33 40 120.0
20240905 60558.693750 60558.707639 0.33 54 162.0
20240913 60566.535417 60566.549306 0.33 78 234.0
20240919 60572.679167 60572.693056 0.33 33 99.0
20240927 60580.632639 60580.646528 0.33 37 111.0
20241005 60588.527778 60588.541667 0.33 34 102.0
20241011 60594.479884 60594.493773 0.33 38 114.0
20241021 60604.520833 60604.534722 0.33 51 153.0
20241029 60612.475000 60612.488889 0.33 41 123.0
20241105 60619.415972 60619.429861 0.33 26 78.0
20241116 60630.405556 60630.488889 2.00 283 141.5
20241123 60637.473611 60637.487500 0.33 47 141.0
20241204 60648.402778 60648.416667 0.33 43 129.0
20241209 60653.353472 60653.436806 2.00 332 166.0
20241216 60660.434028 60660.447917 0.33 31 93.0
20241225 60669.347917 60669.431250 2.00 521 260.5
20241226 60670.399306 60670.413194 0.33 115 345.0
20250101 60676.313889 60676.381944 1.63 653 399.8
20250102 60677.269444 60677.283333 0.33 113 339.0
20250108 60683.306250 60683.347917 1.00 186 186.0
20250112 60687.291667 60687.333333 1.00 475 475.0
20250115 60690.245833 60690.329167 2.00 958 479.0
20250119 60694.281250 60694.302083 0.50 210 420.0
20250123 60698.219444 60698.233333 0.33 79 237.0
20250127 60702.218750 60702.232639 0.33 35 105.0
20250131 60706.240972 60706.282639 1.00 82 82.0
20250208 60714.268056 60714.281944 0.33 17 51.0
20250215 60721.217361 60721.259028 1.00 118 118.0
20250222 60728.248611 60728.262500 0.33 23 69.0
20250301 60735.145139 60735.159028 0.33 20 60.0
20250308 60742.118750 60742.132639 0.33 46 138.0
20250317 60751.089583 60751.103472 0.33 40 120.0
20250322 60756.065278 60756.079167 0.33 39 117.0
20250328 60762.997222 60763.011111 0.33 33 99.0
20250423 60788.921528 60788.935417 0.33 8 24.0
20250501 60796.894444 60796.908333 0.33 57 171.0
20250508 60803.979861 60803.993750 0.33 156 468.0
20250522 60817.883333 60817.897222 0.33 51 153.0
20250525 60820.976389 60820.990278 0.33 49 147.0
20250529 60824.882639 60824.896528 0.33 46 138.0
20250605 60831.968056 60831.981944 0.33 1 3.0
20250612 60838.804167 60838.818056 0.33 0 0.0
20250621 60847.756944 60847.798611 1.00 0 0.0
20250624 60850.750000 60850.770833 0.50 0 0.0
20250713 60869.746528 60869.760417 0.33 0 0.0
20250720 60876.792361 60876.834028 1.00 0 0.0
20250721 60877.711111 60877.731944 0.50 0 0.0
20250723 60879.781250 60879.802083 0.50 0 0.0
20250725 60881.723611 60881.744444 0.50 0 0.0
20250813 60900.753472 60900.767361 0.33 0 0.0
20250824 60911.720139 60911.734028 0.33 0 0.0
20250901 60919.604861 60919.618750 0.33 0 0.0
20250912 60930.589583 60930.603472 0.33 0 0.0
20250921 60939.646528 60939.660417 0.33 0 0.0
20251001 60949.559722 60949.573611 0.33 0 0.0
Table 3: Continued.
Appendix C Search Method for Periodic Signals and QPOs in bursts
@@
To reliably identify periodic signals and quasi-periodic oscillations (QPOs) in the bursts, we adopt the method proposed by [2013ApJ...768...87H], which is based on Bayesian inference. The procedure begins by computing the periodogram of the burst time series using the Fast Fourier Transform (FFT). A simpler model, referred to as the null hypothesis model, is then selected, and the periodogram is fit using maximum likelihood estimation (MLE) to obtain the a maximum posteriori (MAP) parameters under this null hypothesis. The analysis is then performed separately for searching periodic signals and QPOs.
@@
$p\left(\bm{I}|\hat{\theta}_{\rm MAP}^{0},H_{0}\right)=\prod\limits_{i=1}^{n}p\left(I_{j}|S_{j}\right),$ (14)
where $p(I_{j}|S_{j})$ represents the probability density function (PDF) of the periodogram at frequency $f_{j}$ given the model $S_{j}$. Given the model $S_{j}$, the probability density of the power $I_{j}$ is [van1989fourier, 1995A&A...300..707T, 2010MNRAS.402..307V]
$p(I_{j}|S_{j})=\frac{1}{S_{j}}e^{-\frac{I_{j}}{S_{j}}}.$ (15)
The posterior distribution of the null hypothesis model parameters is computed by combining this likelihood with the prior distribution. MCMC sampling is used to obtain a large number of simulated periodograms based on the posterior distribution. The simulated periodograms are then fit with the null hypothesis model, and the maximum residuals for each simulated periodogram are calculated. The $P$-value is determined by comparing the maximum residual observed in the actual burst periodogram with those from the simulated ones. If the $P$-value is below a predefined significance level, the outlier is considered statistically significant and warrants further investigation.
@@
$T_{\rm LRT}^{\rm obs}=-2\log\frac{p\left(\bm{I}|\hat{\theta}_{\rm MAP}^{0},H_{0}\right)}{p\left(\bm{I}|\hat{\theta}_{\rm MAP}^{1},H_{1}\right)}.$ (16)
@@
Our search strategy follows that of [2013ApJ...768...87H] with the following key adjustments: the null hypothesis model consists of a Gaussian model and a constant function model. The Gaussian model fits the low-frequency component of the periodogram, while the constant function model fits the baseline of the periodogram. We found that the Gaussian model provides the best fit for the low-frequency component, and additional Gaussian models are added to fit narrower features in the periodogram. This results in a nested model that serves as the alternative hypothesis. After the initial candidate is identified, the periodogram is reviewed manually to determine if additional Gaussian models are needed. If the periodogram cannot be fit with at most five Gaussian functions plus one constant function (with a reduced $\chi^{2}>1.5$), the candidate is discarded.
@@
Appendix D Supplementary Figures for Quasi-Periodic Signal Analysis